OpenStudy (anonymous):

The bar representing which of the following probabilities on a histogram would be the highest? A.1/30 B.1/15 C.1/10 D.2/15 help asap !! please

jimthompson5910 (jim_thompson5910):

convert each fraction to a decimal number and that will allow you to compare the numbers easier

jimthompson5910 (jim_thompson5910):

for example 1/30 = 0.0333333... 1/15 = 0.0666667 So we can see that 1/15 is larger than 1/30 since 0.0666667 is larger than 0.033333

OpenStudy (anonymous):

so it would be b?

jimthompson5910 (jim_thompson5910):

if you just had those two fractions, yes but you need to compare the rest of the fractions

OpenStudy (anonymous):

c is 0.1 and d is 0.13 so than it would be d than?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

oh okay thank you,

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Do you know anything about this?

jimthompson5910 (jim_thompson5910):

that's the combination of choosing 7 things out of 9 basically it's asking "how many ways are there to choose 7 items out of 9 where order does not matter"

OpenStudy (anonymous):

I don't understand, cause the choices are A.66 B.55 C.36 D.45

jimthompson5910 (jim_thompson5910):

have you learned about permutations?

OpenStudy (anonymous):

no , this is online math,

jimthompson5910 (jim_thompson5910):

I figured that, but I would have also figured that you have learned about permutations or combinations by now

jimthompson5910 (jim_thompson5910):

seems odd how they skipped over that

OpenStudy (anonymous):

No, it was never on any of the information given to me for each section,

jimthompson5910 (jim_thompson5910):

well luckily we can take advantage of the fact that 9 C 7 = 9 C 2 9 C 2 = (9*8)/2 = ???

OpenStudy (anonymous):

36

jimthompson5910 (jim_thompson5910):

so 9 C 7 = 36 as well

jimthompson5910 (jim_thompson5910):

you can also use Pascals triangle to get the same thing

OpenStudy (anonymous):

what about this? If a probability distribution is 5/12,1/6 ,1/4 , x, what is the value of x? A.1/3 B.1/6 C.1/12 D.1/4

jimthompson5910 (jim_thompson5910):

add up all the fractions including x and set it equal to 1 then solve for x

jimthompson5910 (jim_thompson5910):

5/12 + 1/6 + 1/4 + x = 1 solve this for x

OpenStudy (anonymous):

i don't know, 7/22?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

oh wait, one sec

OpenStudy (anonymous):

i don't undersad

OpenStudy (anonymous):

understand*

jimthompson5910 (jim_thompson5910):

what do you get when you combine these fractions 5/12 + 1/6 + 1/4

OpenStudy (anonymous):

i don't know , cause they have different numbers do i need them both to get 12?

jimthompson5910 (jim_thompson5910):

yes you need all the denominators to be equal to 12

OpenStudy (anonymous):

12/12?

OpenStudy (anonymous):

you multiply 6 with 2 and 4 with 3 ,

jimthompson5910 (jim_thompson5910):

5/12 = 5/12 1/6 = 2/12 1/4 = 3/12 ----------------- 5/12 + 1/6 + 1/4 is the same as 5/12 + 2/12 + 3/12

OpenStudy (anonymous):

10/12 reduced to 5/6?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thaat not right

jimthompson5910 (jim_thompson5910):

we're not done yet

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

but we're on the right path to getting there

OpenStudy (anonymous):

okay whats next?

jimthompson5910 (jim_thompson5910):

5/12 + 1/6 + 1/4 + x = 1 is the same as 5/6 + x = 1 Now subtract 5/6 from both sides to isolate x 5/6 + x - 5/6 = 1 - 5/6 x = 1 - 5/6 x = ???

OpenStudy (anonymous):

4/6 reduced to 2/3? but that not on there either?

OpenStudy (anonymous):

nevermind 0.16666666666

jimthompson5910 (jim_thompson5910):

no not this  but this